1.Find the solution of 3x + 6y = 12 when x = 2

2.Given the equation -2y + 4x = 18, find y when x = 1.
3.Write the equation 3x - 4y = 8 in slope-intercept form.
4. The cost Kim pays for chocolates varies directly as the number of pounds of chocolates she buys. She pays $5.40 for 3 pounds of chocolates. Find how much she pays for 5 pounds of chocolates.


Answer by malakumar_kos@yahoo.com(315)

1.Find the solution of 3x + 6y = 12 when x = 2
2.Given the equation -2y + 4x = 18, find y when x = 1.
3.Write the equation 3x - 4y = 8 in slope-intercept form.
4. The cost Kim pays for chocolates varies directly as the number of pounds of chocolates she buys. She pays $5.40 for 3 pounds of chocolates. Find how much she pays for 5 pounds of chocolates.

1) When x= 2 the solution for 3x+6y = 12 will be , 3*2+6y = 12
6+6y = 12
6y = 12-6 = 6
6y = 6
y = 1
The solution for 3x+6y = 12 when x = 2 is y = 1

3) Given eq'n is 3x-4y = 8
this can be expressed in slope intercept form as
-4y = -3x+8 divide by -4 throughout
y = 3x/4-2 is the required slope intercept form

4)Let the cost of chocolates be represented by C
Quantity of chocolates be represented by P
given c is directly proportional to p or C = k*p where k is a
proportionality constant
cost of 3 pounds of chocolates = $5.40 using this we will find
out the value of K
C = K*P
K = C/P = $ 5.40/3 = $ 1.80
therefore cost of 5 pounds of chocolates = C = K*5
= 1.80*5
= 9.00
= $ 9.00



2) -2y+4x = 18 When x = 1 the eq'n becomes -2y = 18-4x

-2y = 18-4*1 = 18-4
-2y = 14 or y = 14/-2
y = -7
The solution for the given eq'n when x = 1 is y = -7